### v1

# A proof of a conjecture of Erdős and Gyárfás on monochromatic path covers

Abstract
In 1995, Erdős and Gyárfás proved that in every $2$-edge-coloured complete graph on $n$ vertices, there exists a collection of $2\sqrt{n}$ monochromatic paths, all of the same colour, which cover the entire vertex set. They conjectured that it is possible to replace $2\sqrt{n}$ by $\sqrt{n}$. We prove this to be true for all sufficiently large $n$.
1 Introduction
In [4], Gerencsér and Gyárfás observed the following fact with a very short and elegant proof.
Theorem 1.1.
The vertex set of any 2-edge-coloured complete graph $K_{n}$ can be covered by two monochromatic paths.
Importantly, the two monochromatic paths do not have to be of the same colour. In his survey [7], Gyárfás mentions that when he first told Erdős about this result, Erdős misunderstood this detail, which led the two of them to investigate the smallest number of monochromatic paths of the same colour that cover all vertices of a 2-edge-coloured complete graph. In [2], they proved the following.
Theorem 1.2.
In every $2$-edge-coloured complete graph on $n$ vertices, there exists a collection of $2\sqrt{n}$ monochromatic paths, all of the same colour, which cover the entire vertex set.
The bound $2\sqrt{n}$ in Theorem 1.2 is best possible up to a constant by the following construction. Suppose $\sqrt{n}\in\mathbb{N}$ and partition $V(K_{n})$ into disjoint sets $A$ and $B$ of order $n-\sqrt{n}+1$ and $\sqrt{n}-1$ respectively. Colour the edges contained in $A$ blue, and all other edges red. Each red path covers at most $\sqrt{n}$ vertices in $A$ since it must alternate between $A$ and $B$, so we need at least $\lceil|A|/\sqrt{n}\rceil=\sqrt{n}$ red paths to cover the entirety of $V(K_{n})$. Every collection of blue paths covering $V(K_{n})$ must include all individual vertices in $B$ as distinct paths and needs another blue path to cover $A$, thus has size at least $|B|+1=\sqrt{n}$. If $n$ is not a perfect square, one can use the same construction with $|B|=\lfloor\sqrt{n}\rfloor-1$, so that any cover requires at least $\lfloor\sqrt{n}\rfloor$ paths of the same colour.
Erdős and Gyárfás conjectured (see [7]) that this construction is the colouring that requires the most paths for a cover, meaning that the bound in their theorem could be improved to $\sqrt{n}$. In this paper, we confirm their conjecture for all sufficiently large $n$.
Theorem 1.3.
There exists $n_{0}\in\mathbb{N}$ such that for all $n>n_{0}$ and all 2-edge-colourings of $E(K_{n})$, there exists a collection of at most $\sqrt{n}$ monochromatic paths of the same colour that cover the vertices of $K_{n}$.
The constant $n_{0}$ that we find is large, but with some additional technical effort, one can show that for all $n\in\mathbb{N}$, $\sqrt{n}+10$ monochromatic paths of the same colour are sufficient to cover $K_{n}$.
Several related problems of this type have also been studied, looking for monochromatic structures that either cover or partition the vertex set within an edge-coloured graph. For example, Lehel’s conjecture, proven in full by Bessy and Thomassé [1], strengthens Theorem 1.1. It states that the vertices of a $2$-edge-coloured $K_{n}$ can be covered by two vertex-disjoint monochromatic cycles (of course, they must be allowed to have different colours). A famous generalisation of Lehel’s conjecture due to Gyárfás [5], stating that an $r$-edge-coloured complete graph can be covered by $r$ monochromatic cycles, remains wide open. We refer the reader to the survey [7] for partial results and further related problems.
Finally, as an interpolation between Theorem 1.1 and Theorem 1.2, Eugster and Mousset [3] investigated the minimal number of monochromatic paths needed to cover the $r$-edge-coloured $K_{n}$ if at most $s\leq r$ different colours are allowed to be used for the paths in total.
2 Preliminaries
We call a collection $\mathcal{P}$ of paths in a red-blue edge-coloured graph $G$ a red path cover if every vertex in $V(G)$ is covered by some path in $\mathcal{P}$, and every $P\in\mathcal{P}$ contains only red edges. We define a blue path cover similarly. The length of a path is the number of edges it contains, and importantly we allow paths to have length zero, so that a path can be a single vertex. Note that paths need not be disjoint, unlike those required for some definitions of the path covering number of a graph.
We will use a bipartite Ramsey-type result for paths determined by Gyárfás and Lehel in [6] (see Theorem 3 and Remark 1).
Lemma 2.1.
If $k$ and $\ell$ are distinct natural numbers then every red-blue edge-coloured complete bipartite graph $K_{\left\lceil\frac{k+l}{2}\right\rceil,\left\lceil\frac{k+l}{2}\right\rceil}$ contains a red path of length $k$ or a blue path of length $\ell$.
Suppose we have found a long monochromatic path $P$ in a red-blue edge-coloured $K_{n}$, without loss of generality we may assume $P$ is blue. We may consider the bipartite graph between $V(P)$ and a subset of $V(K_{n})\setminus V(P)$, comprised only of red edges. If there is a long red path, say $R$, is this bipartition then we obtain a large set $V(P)\cap V(R)$ covered by both a red path and a blue path, which allows us some flexibility for covering the remainder of the vertices. A similar idea may be used if we can find few red paths that cover many vertices in the bipartition. In the following two lemmas we consider certain degree properties within bipartite graphs that allow us to find few paths covering many vertices.
Lemma 2.2.
Let $G$ be a bipartite graph on vertex set $X\cup Y$ with the property that every vertex in $Y$ has degree at least $(|X|+|Y|)/2$. Then $G$ contains a path covering $2|Y|$ vertices.
Proof.
Let $P=x_{1}y_{1}\ldots x_{k}y_{k}$ be a path starting in $X$ and ending in $Y$ and suppose that $k<|Y|$. Fix $y_{k+1}\in Y\setminus\{y_{1},\ldots,y_{k}\}$. We have
$|N(y_{k})\cap N(y_{k+1})|\geq\frac{|X|+|Y|}{2}-\left(|X|-\frac{|X|+|Y|}{2}\right)=|Y|>k.$
Therefore, $y_{k}$ and $y_{k+1}$ have a common neighbour $x_{k+1}\in X\setminus\{x_{1},\ldots,x_{k}\}$, and we can extend $P$ via $y_{k}x_{k+1}y_{k+1}$. ∎
Lemma 2.3.
Let $m\in\mathbb{N}$ and let $G$ be a bipartite graph on vertex set $X\cup Y$ with $|X|\geq|Y|+2m$, and such that every vertex in $Y$ has degree at least $|X|-m$. Then $G$ contains a collection of at most $\lfloor|X|/|Y|\rfloor$ paths whose union covers all but $|Y|+2m$ vertices in $X$ and all vertices in $Y$.
Proof.
We prove the statement by induction on $|X|$, always assuming that $|X|\geq|Y|+2m$. Since
$|X|-m\geq\frac{|X|+|Y|}{2},$
we may apply Lemma 2.2, to obtain a path $P$ covering $2|Y|$ vertices in $G$. In particular, $P$ covers all vertices in $Y$. Let $X^{\prime}=X\setminus V(P)$, and note that in the induced subgraph $G^{\prime}=G[X^{\prime}\cup Y]$, every vertex in $Y$ has degree at least $|X^{\prime}|-m$. If $|X^{\prime}|\leq|Y|+2m$, then $\{P\}$ is a collection of paths as desired. Otherwise, we may apply the induction hypothesis to $G[X^{\prime}\cup Y]$ to obtain a collection of paths $P_{1},\ldots,P_{k}$ with $k\leq\lfloor|X^{\prime}|/|Y|\rfloor=\lfloor|X|/|Y|\rfloor-1$ that cover all but $|Y|+2m$ vertices in $X^{\prime}$. Adding $P$ to this collection yields the claim. ∎
To get the exact bound in Theorem 1.3, we also require a version of Lemma 2.3 which lets us cover all vertices in $X$ when many vertices in $Y$ are adjacent to all vertices in $X$.
Lemma 2.4.
Let $G$ be a non-empty bipartite graph on vertex set $X\cup Y$ with sets $X_{0}=\{x\in X:d(x)=|Y|\}$, $X_{1}=X\setminus X_{0}$, $Y_{0}=\{y\in Y:d(y)=|X|\}$, and $Y_{1}=Y\setminus Y_{0}$ such that
  1. (i)
    $|X|>|Y|$, and
  2. (ii)
    $X_{1}=Y_{1}=\emptyset$ or $|X_{0}|/|Y_{1}|>2|X_{1}|/|Y_{0}|$.
Then $G$ contains a collection of at most $\left\lceil\frac{|X|}{|Y|+1}\right\rceil$ paths that cover all vertices in $G$.
Proof.
We prove the statement again by induction on $|X|$. It follows from (i) and (ii) that $|X_{0}|\geq|Y_{1}|+1$, and therefore, $G$ contains a path $P$ with $12|Y_{1}|+$ 1 vertices, alternating between $X_{0}$ and $Y_{1}$, such that $P$ starts at $x_{0}\in X_{0}$ and also ends in $X_{0}$. In particular $P$ covers all vertices in $Y_{1}$.
Furthermore, since every vertex in $Y_{0}$ has neighbourhood $X$, and $|X|>|Y|$, $G$ contains a path $Q$ with $2\min\{|Y_{0}|,|X_{1}|\}$ vertices, alternating between $Y_{0}$ and $X\setminus V(P)$, that starts at $y_{0}\in Y_{0}$ and covers $\min\{|Y_{0}|,|X_{1}|\}$ vertices in $X_{1}$. Consider the concatenated path $R=Px_{0}y_{0}Q$ and the graph $G^{\prime}=G[X^{\prime}\cup Y]$ where $X^{\prime}=X\setminus V(R)$.
Defining the sets $X_{0}^{\prime},X_{1}^{\prime},Y_{0}^{\prime}$, and $Y_{1}^{\prime}$ analogously for $G^{\prime}$ to $X_{0},X_{1},Y_{0}$, and $Y_{1}$ were defined for $G$ respectively, we observe that $Y_{1}^{\prime}\subset Y_{1}$ and $X_{1}^{\prime}\subset X_{1}\setminus V(Q)$, while $Y_{0}\subset Y_{0}^{\prime}$. If we assume that $X_{1}^{\prime}=\emptyset$ or $Y_{1}^{\prime}=\emptyset$, then it follows that $G^{\prime}$ is complete bipartite and it is trivially possible to cover $X\setminus V(R)$ with $\left\lceil\frac{|X^{\prime}|}{|Y|+1}\right\rceil=\left\lceil\frac{|X|}{|Y|+1}\right\rceil-1$ paths so that nothing is left to show. Suppose therefore that $X_{1},Y_{1}\neq\emptyset$, in particular, we must have $|X_{1}|>|Y_{0}|$ so that $|V(Q)\cap X_{1}|=|Y_{0}|$. Using this, we see that
$\frac{|X_{0}^{\prime}|}{|Y_{1}^{\prime}|}\geq\frac{|X_{0}\setminus V(P)|}{|Y_{1}|}\geq\frac{|X_{0}|}{|Y_{1}|}-2>2\left(\frac{|X_{1}|}{|Y_{0}|}-1\right)=\frac{2|X_{1}\setminus V(Q)|}{|Y_{0}|}>\frac{2|X_{1}^{\prime}|}{|Y_{0}^{\prime}|}.$
If $|X^{\prime}|>|Y^{\prime}|$, we may use the induction hypothesis to obtain a family $\mathcal{P}$ of at most $\left\lceil\frac{|X|}{|Y|+1}\right\rceil-1$ paths which cover all of $X^{\prime}$ and $Y$, and together with $R$, these paths cover all of $G$.
If $|X^{\prime}|\leq|Y^{\prime}|$ on the other hand, then we must have $|Y_{0}^{\prime}|>|X_{1}^{\prime}|$ so that it is possible to cover all of $X_{1}^{\prime}$ with a single path $Q^{\prime}$ of length $2|X_{1}^{\prime}|$. The remaining graph $G^{\prime}-V(Q^{\prime})$ is complete bipartite so that we may find a path $P^{\prime}$ that covers all of $X_{0}^{\prime}$ in it. Concatenating $P^{\prime}$ and $Q^{\prime}$ suitably, we obtain a second path $R^{\prime}$ such that $R$ and $R^{\prime}$ cover all of $G$. ∎
3 Proof of Theorem 1.3
Let $n\in\mathbb{N}$. In what follows, without loss of generality we assume the vertex set of $K_{n}$ is given by $[n]$. Let $\mathcal{X}_{n}$ be the set of all red-blue colourings of $E(K_{n})$. For $\chi\in\mathcal{X}_{n}$, denote by $f(n,\chi)$ the smallest family of monochromatic paths of the same colour that cover $[n]$. Furthermore, let $f(n)=\max_{\chi\in\mathcal{X}_{n}}f(n,\chi)$.
In order to prove Theorem 1.3, we follow an inductive strategy to initially prove a weaker upper bound on $f(n)$ for all $n\in\mathbb{N}$, given in Proposition 3.4 which can then be bootstrapped to prove our main theorem. Thus in each of the lemmas in this section, we will assume an upper bound on $f(m)$ for all $m<n$ that will relate to our inductive hypothesis.
We first introduce a lemma which tells us that under a colouring $\chi\in\mathcal{X}_{n}$, if there exists a large set $S$ that is covered by both a collection of few red paths, and a collection of few blue paths, then we can obtain a strong upper bound on $f(n,\chi)$.
Lemma 3.1.
Let $C_{1}\geq C_{2}\geq 0$ and let $n\in\mathbb{N}$ be such that $f(m)<\sqrt{m}+C_{1}$ for all $m<n$. Let further $k\in\mathbb{N}$ and $\chi\in\mathcal{X}_{n}$ be such that there exist red paths $P_{1},\ldots,P_{k}\subset K_{n}$, blue paths $Q_{1},\ldots,Q_{k}\subset K_{n}$, and a set $S\subset[n]$ satisfying $S\subset(V(P_{1})\cup\ldots\cup V(P_{k}))\cap(V(Q_{1})\cup\ldots\cup V(Q_{k}))$. If
$\sqrt{n-|S|}+C_{1}+k\leq\sqrt{n}+C_{2},$
then $f(n,\chi)\leq\sqrt{n}+C_{2}$. In particular, if there is a red path $P$ and a blue path $Q$ such that $S\subset V(P)\cap V(Q)$, and
$|S|\geq 2(C_{1}-C_{2}+1)\sqrt{n},$
then $f(n,\chi)\leq\sqrt{n}+C_{2}$.
Proof.
Suppose there exist red paths $P_{1},\ldots,P_{k}\subset K_{n}$, blue paths $Q_{1},\ldots,Q_{k}\subset K_{n}$, and a set $S\subset[n]$ satisfying $S\subset(V(P_{1})\cup\ldots\cup V(P_{k}))\cap(V(Q_{1})\cup\ldots\cup V(Q_{k}))$. Let $H$ be the 2-edge-coloured graph obtained by deleting $S$ from $K_{n}$ under the colouring $\chi$. Then $H$ has $m=n-|S|$ vertices, and by definition of $f(m)$, there exists a collection of at most $f(m)$ monochromatic paths of the same colour which cover $V(H)$. Adding $\{P_{1},\ldots,P_{k}\}$ or $\{Q_{1},\ldots,Q_{k}\}$ to this collection according to whether these paths are red or blue respectively yields a collection of at most $f(m)+k$ monochromatic paths of the same colour covering $V(H)\cup S=[n]$ in $K_{n}$. By our assumptions, we have
$f(n,\chi)\leq f(m)+k<\sqrt{m}+C_{1}+k=\sqrt{n-|S|}+C_{1}+k\leq\sqrt{n}+C_{2},$
as desired. For the second part of the statement, observe that if there is a red path $P$ and a blue path $Q$ such that $S\subset V(P)\cap V(Q)$ for some $S\subset[n]$ with $|S|\geq 2(C_{1}-C_{2}+1)\sqrt{n}$, then
$\sqrt{n-|S|}\leq\sqrt{n-2(C_{1}-C_{2}+1)}\leq\sqrt{n}-(C_{1}-C_{2}+1).$
Rearranging gives
$\sqrt{n-|S|}+C_{1}+1\leq\sqrt{n}+C_{2},$
and so taking $k=1$, we deduce that $f(n,\chi)\leq\sqrt{n}+C_{2}$. ∎
We now show that if $\chi\in\mathcal{X}_{n}$ is such that there exists no small monochromatic path cover, then there must exist some long monochromatic path, of colour $\gamma$ say, such that for every vertex $y$ outside of the path, there are few vertices $x$ on the path such that $\chi(xy)=\gamma$.
Lemma 3.2.
Let $C_{1}\geq C_{2}\geq 0$ and let $n>8^{4}(C_{1}-C_{2}+1)^{4}$ be such that $f(m)<\sqrt{m}+C_{1}$ for all $m<n$. If $\chi\in\mathcal{X}_{n}$ is such that $f(n,\chi)>\sqrt{n}+C_{2}$, then there a monochromatic path $P$ such that the size of $Y=[n]\setminus V(P)$ is at most $\sqrt{n}+8(C_{1}-C_{2}+1)\sqrt[4]{n}$. Furthermore, if $\gamma$ is the colour of the edges of $P$, then for every $y\in Y$, the size of $\{x\in V(P):\chi(xy)=\gamma\}$ is at most $2(C_{1}-C_{2}+1)\sqrt{n}$.
Proof.
Let $\Delta=C_{1}-C_{2}$ and let $\chi\in\mathcal{X}_{n}$ be such that $f(n,\chi)>\sqrt{n}+C_{2}$. Without loss of generality, there exists a blue path $Q$ which covers $\lfloor n/2\rfloor$ vertices. Let $W$ be a subset of $[n]\setminus V(Q)$ of size $\lfloor n/2\rfloor$ and consider the longest red path $R$ in the complete bipartite graph between $V(Q)$ and $W$. The set $V(R)\cap V(Q)$ is covered both by a single red and a single blue path so that by Lemma 3.1, we may assume that $|V(R)\cap V(Q)|<2(\Delta+1)\sqrt{n}$. However, since $R$ is a path in a bipartite graph between $V(Q)$ and $W$, at least $\lfloor|V(R)|/2\rfloor$ vertices of $R$ must lie in $V(Q)$, and thus, it follows that $|V(R)|<4(\Delta+1)\sqrt{n}$. By Lemma 2.1, there exists a blue path of length at least $n-2(\Delta+1)\sqrt{n}-1$ between $V(Q)$ and $W$.
Let $P$ be the longest blue path in $K_{n}$, denote its vertex set by $X$, and let $Y=[n]\setminus X$. Consider $y\in Y$ and let $B$ denote the set of blue neighbours of $y$ in $P$. Since $P$ cannot be extended, $B$ does not contain any endpoint or any two consecutive vertices of $P$. Furthermore, if we direct $P$ arbitrarily, the predecessors of $B$ form a red clique set. Indeed, if $P=v_{1}v_{2}\ldots v_{k}$, $1\leq i<j\leq k$, and $v_{i}v_{j}$, $v_{i+1}y$, and $v_{j+1}y$ were all blue, then $v_{1}\ldots v_{i}v_{j}v_{j-1}\ldots v_{i+1}yv_{j+1}\ldots v_{k}$ would be a blue path of length $k+1$. However, we may argue as before that $V(P)$ cannot contain a red path of length $2(\Delta+1)\sqrt{n}$, which means that $y$ has at most $2(\Delta+1)\sqrt{n}$ blue neighbours in $X$.
In other words, taking $m=2(\Delta+1)\sqrt{n}$, each vertex in $Y$ has at least $|X|-m$ blue neighbours in $X$. What is more, we know that $|X|\geq n-2(\Delta+1)\sqrt{n}-1$ and hence $|Y|\leq 2(\Delta+1)\sqrt{n}+1$. Thus, as $n>100(\Delta+1)^{2}$,
$|X|\geq n-2(\Delta+1)\sqrt{n}-1>10(\Delta+1)\sqrt{n}-2(\Delta+1)\sqrt{n}-1>6(\Delta+1)\sqrt{n}+1\geq|Y|+2m.$
It now follows from Lemma 2.3 that the bipartite graph given by the red edges between $X$ and $Y$ contains a collection of at most $\lfloor|X|/|Y|\rfloor<n/|Y|$ paths which cover all but $|Y|+2m<6(\Delta+1)\sqrt{n}+1$ vertices in $X$ and all vertices of $Y$.
However, all vertices in $X$ that are covered by this collection of paths are also covered by the single blue path $P$, which means that by Lemma 3.1,
$\sqrt{n}+C_{2}<\sqrt{6(\Delta+1)\sqrt{n}+1}+C_{1}+n/|Y|\leq 3(\Delta+1)\sqrt[4]{n}+C_{1}+n/|Y|,$
which we can simplify to
$\sqrt{n}<n/|Y|+4(\Delta+1)\sqrt[4]{n},$
which implies that
$|Y|<\frac{n}{\sqrt{n}-4(\Delta+1)\sqrt[4]{n}}=\frac{\sqrt{n}}{1-4(\Delta+1)/\sqrt[4]{n}}.$
Since we have $4]{n}\geq 8(\Delta+1)$ by assumption and $1/(1-x)<1+2x$ for $x\in[0,1/2]$, we get that
$|Y|<\sqrt{n}+8(\Delta+1)\sqrt[4]{n},$
as desired. ∎
If there exists a monochromatic path in $K_{n}$ containing almost all vertices, we can afford to cover the remaining vertices individually or in pairs. The following lemma formalises this.
Lemma 3.3.
Let $n\in\mathbb{N}$ and $\chi\in\mathcal{X}_{n}$ and suppose $P$ is a blue path in $K_{n}$. Let $Y=[n]\setminus V(P)$ and write $Y_{0}$ for the set of vertices in $Y$ that are not contained in a blue edge with any vertex in $V(P)$. Then we have
$f(n,\chi)\leq 1+\left\lceil\frac{|Y\setminus Y_{0}|}{2}\right\rceil+|Y_{0}|<2+|Y|/2+|Y_{0}|/2.$
Proof.
Let $Y_{1}\coloneqq Y\setminus Y_{0}$ and note that for any two vertices $y_{1},y_{2}\in Y_{1}$, there exists a blue path between $y_{1}$ and $y_{2}$ whose interior points lie in $P$. Thus, by partitioning the vertices in $Y_{1}$ arbitrarily into pairs, we may cover all of them with at most $\lceil|Y_{1}|/2\rceil$ blue paths. Adding a single-vertex path for each $y\in Y_{0}$ and $P$ to this collection, we obtain a blue path cover of $K_{n}$ of size at most $1+\lceil|Y_{1}|/2\rceil+|Y_{0}|$. Since any monochromatic path cover of $[n]$ must have size at least $f(n,\chi)$, this yields
$f(n,\chi)\leq 1+\left\lceil\frac{|Y_{1}|}{2}\right\rceil+|Y_{0}|<2+\frac{|Y_{1}|}{2}+|Y_{0}|=2+\frac{|Y|-|Y_{0}|}{2}+|Y_{0}|=2+\frac{|Y|}{2}+\frac{|Y_{0}|}{2}.$
We now prove a weak version of Theorem 1.3, which will bootstrap to prove Theorem 1.3 in full.
Proposition 3.4.
For all $n\in\mathbb{N}$ and $C=160000$, $f(n)<\sqrt{n}+C$.
Proof of Proposition 3.4.
We prove the claim by induction on $n$, noting that it is trivial for $n\leq C$. Suppose therefore that $n>C$ and that $f(m)<\sqrt{m}+C$ for all $m<n$ and consider a colouring $\chi\in\mathcal{X}_{n}$. Applying Lemma 3.2 with $C=C_{1}=C_{2}$, we obtain without loss of generality a blue path $P$ such that the size of $Y=[n]\setminus V(P)$ is at most $\sqrt{n}+8\sqrt[4]{n}<3\sqrt{n}/2$ and such that every $y\in Y$ has at least $|X|-2\sqrt{n}$ blue neighbours in $X=V(P)$. By Lemma 3.3, if $|Y_{0}|\leq\sqrt{n}/2$ or $|Y|\leq\sqrt{n}$, then $f(n,\chi)\leq\sqrt{n}+2$ and there is nothing left to show. Let us therefore assume that $|Y_{0}|>\sqrt{n}/2$ and $|Y|>\sqrt{n}$.
On the other hand, we have that $|X|\geq n-3\sqrt{n}/2\geq 3\sqrt{n}/2+2m$ for $m=2\sqrt{n}$. Therefore, by applying Lemma 2.3, we see that there must exist a collection of at most $\sqrt{n}$ red paths that cover all vertices in $Y$ and all but at most $6\sqrt{n}$ vertices in $X$. Let $X^{\prime}\subset X$ denote the vertices that are not covered by these paths and recall that all edges between $Y_{0}$ and $X$, in particular those between $Y_{0}$ and $X^{\prime}$, are red. Since $|Y_{0}|>\sqrt{n}/2$ and $|X^{\prime}|<6\sqrt{n}$ we can cover all of $X^{\prime}$ with at most 12 red paths between $X^{\prime}$ and $Y_{0}$, which completes the proof. ∎
We are now ready to prove Theorem 1.3.
Proof of Theorem 1.3..
Let $C=160000$, $n_{0}=C^{10}$ and $n>n_{0}$. Define $\alpha_{n}=9C/\sqrt[4]{n}$, and suppose $\chi\in\mathcal{X}_{n}$ is such that $f(n,\chi)>\sqrt{n}$. By Proposition 3.4, we have $f(m)<\sqrt{m}+C$ for every $m\in\mathbb{N}$. Applying Lemma 3.2 with $C_{1}=C$ and $C_{2}=0$, we see that without loss of generality, there exists a blue path $P$ such that the size of $Y=[n]\setminus V(P)$ is at most $(1+\alpha_{n})\sqrt{n}$ and with the additional property that every vertex in $Y$ has less than $2C\sqrt{n}$ blue neighbours in $X=V(P)$. By Lemma 3.3, if $|Y_{0}|\leq(1-2\alpha_{n})\sqrt{n}$ or $|Y|\leq\sqrt{n}-1$, then $f(n,\chi)\leq\sqrt{n}$, a contradiction. Let us therefore assume that $|Y_{0}|>(1-2\alpha_{n})\sqrt{n}$ and $|Y|>\sqrt{n}-1$.
We define $H$ as the bipartite graph $H$ on vertex set $X\cup Y$ whose edges are the red edges between $X$ and $Y$ under $\chi$. Let $Y_{1}=Y\setminus Y_{0}$, $X_{0}=\{x\in X:d_{H}(x)=|Y|\}$, and $X_{1}=X\setminus X_{0}$. We have $|Y_{1}|=|Y|-|Y_{0}|<(1+\alpha_{n})\sqrt{n}-(1-2\alpha_{n})\sqrt{n}=27Cn^{1/4}$. By choice of $P$, we know that |$|\{x\in X:\chi(xy)=\text{blue}\}|<2C\sqrt{n}$ for every $y\in Y$, which implies that $|X_{1}|<2C\sqrt{n}|Y_{1}|<54C^{2}n^{3/4}$. Now, either $Y_{1}=\emptyset=X_{1}$, or, since $n$ is sufficiently large, we have
$\frac{|X_{0}|}{|Y_{1}|}=\frac{|X|-|X_{1}|}{|Y_{1}|}>\frac{n-(1+\alpha_{n})n^{1/2}-54C^{2}n^{3/4}}{27Cn^{1/4}}>n^{1/2}>\frac{108C^{2}n^{3/4}}{(1-2\alpha_{n})n^{1/2}}>\frac{2|X_{1}|}{|Y_{0}|}.$
Either way, we may apply Lemma 2.4 to obtain a collection of at most $\lceil|X|/(|Y|+1)\rceil$ paths in $H$ covering all vertices in $H$. This corresponds to a red path cover in $K_{n}$, which implies that $f(n,\chi)\leq\lceil|X|/(|Y|+1)\rceil$. Since $|Y|>\sqrt{n}-1$, it follows that
$f(n,\chi)\leq\left\lceil\frac{|X|}{|Y|+1}\right\rceil=\left\lceil\frac{n+1}{|Y|+1}\right\rceil-1\leq\left\lceil\sqrt{n+1}\right\rceil-1\leq\sqrt{n}.$
Thus we have a contradiction, proving the theorem. ∎
Acknowledgements
The second author is supported by the London Mathematical Society and the Heilbronn Institute for Mathematical Research through an Early Career Fellowship. The third author is supported by the Martingale Foundation.